The ΔG°\' of the reaction is -6.220 kJ ·mol–1. Calculate the equilibrium constant for the reaction. (Assume a temperature of 25° C.)
Answer #1

T= 25.0 oC

= (25.0+273) K

= 298 K

ΔGo = -6.22 KJ/mol

ΔGo = -6220 J/mol

use:

ΔGo = -R*T*ln Kc

-6220 = - 8.314*298.0* ln(Kc)

ln Kc = 2.5105

Kc = 12.31

Answer: 12.31